A cargo plain flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220 mph. The average speen on the way back was 200 mph. How many hours did the trip there take?
Answer by mananth(16949)

A cargo plain flew to the maintenance facility and back. It took one hour less time to get there than it did to get back. The average speed on the trip there was 220 mph. The average speen on the way back was 200 mph. How many hours did the trip there take?
Let the plane take x hours to return.
it took x-1 hours to reach there.
Total time taken for the trip = x+x-1 = 2x-1
Total distance = total time * average speed
2x-1 * 220 = total distance
.
speed on return = 220 mph
time = x hours
200x is the distance on return. Since the total distance is double of one way distance
400x is the total distance
220(x-1) =400x
440x-220 = 400x
440x-400x = 220
40x=220
x= 5.4 hours time to return.
.
time to go = x-1
= x-1
=4.4 hours
Total time taken = 9.8 hours.